Consider a B-G cavity chamber with equilibrium thickness copper walls. It is first filled with a mass m of air, then by the same mass of hydrogen. Assumingidentical y-ray irradiations in the two cases, what is the charge ratioAssume ionic recombination has been corrected for, that (w le)H= 36.5JIC, and that the mean electron energy = 0.80 MeV.
Answer #1

given ha the nean electsnon enesoev 0 er T7 att J leSa H e ohasige of teio Assure ionic
Answer #2

Screenshot_1.jpg

Excuse me, I have a question How do you get the term that I marked in red in the photo? please.

answered by: Fabiola
Answer #3

D = (q/m)(W/e)


q = D*m/(W/e)


(qair/qH) = (Dair/DH)*(mair/mH)*[(W/e)H/(W/e)air]


*mair=mH

*(W/e)air = 33.97 J/C


(qair/qH) = (1.683/3.883)*(36.5/33.97)


(qair/qH) = 0.466